What's the empirical formula of a molecule containing 65.5% carbon, 5.5% hydrogen and 29.0% oxygen?
1
Expert's answer
2011-05-17T11:32:15-0400
C(65.5%), H(5.5%), O(29%)
Molar weights of the elements are:
m(C) = 12, m(H) = 1, m(O) = 16.
65.5 : 5.5 : 29 = 12 x : 1 y : 16 z

65.5/5.5 = 12x/y, x/y = 0.99 x = y
29/5.5 = 16z/y, z/y = 0.33 z = 1/3 y
65/29 = 12x / 16z, x/z = 3 x = 3z
Thus we can find that if z = 1, x = y = 3, the empirical formula would be C3H3O.